A 1 m diameter sphere is droped from an evacuated 145 m drop tower.
1) how long is sphere in free fall ??
2) when it reaches bottom, it experiences an avg decceleration of 25 g  . Through wat dist it travel during decceleration ??

Grade:12

1 Answers

Gokul Joshi AskiitiansExpert-IITK
13 years ago

Dear Mohit,

s = ut + 1/2 at2 (where a = g)
145 = 0 + 1/2 10 t2
t = sqrt(29) s

Second question is a bit shady. From average decceleration one can't calculate the distance, only change in velocity. More precisely it should have mentioned "CONSTANT" decceleration. Then

v2 = u2 + 2as will give you the answer where u = at = 10*sqrt(29) and v = 0.


Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly. All the best.

Regards,
Askiitians Experts,
Gokul  Joshi